OpenStudy (anonymous):

a conducting loop is painted around the equator of a spherical rubber balloon.a magnetic field B=0.2cos4t is applied perpendicular to the plane of the equator.the balloon is contracting inwardly with a radial velocity v.when the balloon radius is r=0.5m,the rms voltage induced in the loop is 500mV.find the velocity v at this instant? answer:0.52m/s

OpenStudy (anonymous):

E.M.F = -D/Dt( Flux ) Now, Magnetic Flux F = B.A , both B & A vary with time Here B = BoCos(wt) where Bo = 0.2, w = 4 also, dr/dt = -v , radially inwards Hence, V = -D/Dt( B.A) = -d/dt(A* d(B)/Dt + B * dA/dt) = -D/Dt( BoCostwt * pi * r^2) (we get 2 terms on differentiating) = ( -wBoSin(wt). pi*r^2 - BoCoswt*2*pi*r*v) V = pi*r*Bo (2v Coswt + wr Sinwt) lets convert this to Cos(A-B) form using CosACosB + SinASinB formula V = \[ \pi*r*Bo \sqrt{(rw)^2 + (2v)^2} * \cos (wt - \delta)\] where \[\tan \delta = (r*w/2v)\] Thus we get our induced voltage, \[V = Vo \cos (wt - \delta)\] where \[Vo = \pi*r*Bo*\sqrt{(r*w)^2 + (2*v)^2}\] Now, in our question, V(R.M.S) is given = 0.5 Volt \[Vrms = Vo/\sqrt{2}\] Thefore - \[0.5 = \pi*r*Bo*\sqrt{(r*w)^2 + (2*v)^2}/\sqrt{2}\] solve for v, everything else is known Bo = 0.2, r = 0.5 , w = 4 we get v = 0.5m/s

OpenStudy (anonymous):

I have also attached this file which has my original hand written solution. I couldn't write it properly above using the equations editor on this page

OpenStudy (anonymous):

i found it by other formula since in this pblm we have motional emf we take r=vt. the formula for emf is e=v*b*l e=v*b*2*pi*r(since r=vt) e=v^2*b*2*pi*t from this velocity,v=sqrt(e/2*pi*b*t)------(1) omega,w=2*pi*f=4(given) f=2/pi; t=pi/2=1.57s e=.5v,b=.2 T, substituting these values in(1) we get v=0.51m/s. anyways,your approach was also very nice

OpenStudy (anonymous):

Hey @ananta, thats a nice shortcut you thought of.. one thing I'm wondering about though.. you calculated t(time elapsed until now when radius is 0.5 and velocity = v ) by equating it to the time period of the magnetic field function i.e. t = 1/f = Time period. What is your reasoning for that ? How would you use this formula, if say, the given radius and consequent EMF were given for a different instant, which would mean the time elapsed would be different and you wouldn't be able to equate it to Magentic Field function's time period ?